### v1

# Stochastic interpretation of quantum mechanics

Abstract
We express the probabilistic character associated to the wave function by treating it as a stochastic variable. This is accomplished by means of a stochastic equation for the wave function whose noise changes the phase of the wave function but not its absolute value, so that the norm of the wave function is strictly conserved along a stochastic trajectory. We show that the density matrix that obeys the quantum Liouville equation is the covariance matrix associated to the stochastic wave function.
I Introduction
Quantum mechanics landau1958 ; merzbacher1961 ; messiah1961 ; sakurai1967 ; sakurai1994 ; griffiths1995 ; piza2002 ; griffiths2002 differs fundamentally from classical mechanics lanczos1949 ; goldstein1950 ; landau1960 ; arnold1978 . In contrast to the latter, the former contains quantities that do not have obvious interpretation. In classical mechanics the state of a system is defined by positions an momenta of the particles, which are interpreted as the real positions and the real momenta of the particles. In quantum mechanics, on the other hand, the state is defined by the wave function, which has no obvious interpretation.
The wave function is not observed experimentally and for that reason it is called unobservable. For many, this terminology implies, usually in a tacit manner, a metaphysical conception of the wave function. That is, a concept that is real but which is over and beyond all experience carnap1935 . However, a reality which is beyond experience is inconsistent and contradictory, at least in the realm of science. If we deny such an understanding we may be lead to propositions that are devoid of scientific meaning or even unscientific. In fact, there is no need to assign a real existence of an unobservable if we understand a scientific theory as a representation of the real and not the real itself.
In accordance with the standard interpretation of quantum mechanics omnes1994 ; auletta2001 ; barandes2020 ; freire2022 ; barioni2022 ; pessoa2025 the square of the absolute value of the wave function is related to probability. More precisely, if $\psi(x)$ is the wave function associate to the position $x$ of a particle that moves along a straight line then $|\psi(x)|dx$ is the probability of finding the particle between $x$ and $x+dx$. However, we need also to state the meaning or the interpretation of probability. Within the standard interpretation, it seems that it is interpretated as related to frequency, which can be measured experimentally.
In accordance with the standard interpretation, the probabilistic character of the wave function emerges at the instant of measurement, when the wave function is said to collapse. A question then arises as to the probabilistic character of the wave function prior to measurement. That is, suppose that we perform a measurement at time $t$ and ask whether at a time $t^{\prime}<t$ the wave function has the probabilistic character. To answer this question we must take a measurement at the instant $t^{\prime}$. But, if we do that the wave function collapses at $t$ and the measurement at time $t$ becomes compromised.
From these reasonings, we may say that the probabilistic character of the wave function prior to measurement is not li ble to measurement. Within the standard interpretation of quantum mechanics, it is assumed that the wave function is not associated to probability prior to measurement, and that the probabilistic character emerge at the instant of measurement. Our aim here is to provide an approach in which the wave function does have a probabilistic character at any time. This is acomplished by treating the wave function as a stochastic variable which obeys a stochastic equation of motion oliveira2024a ; oliveira2024b ; oliveira2025 . It is important to say that the probability character at the instant of measurement given by the present approach is the same as that given by the standard interpretation.
II Stochastic equations of motion
We consider a state vector $|\psi\rangle$ defined on a Hilbert space of dimension $n$,
$|\psi\rangle=\sum_{j}z_{j}|j\rangle,$ (1)
where $z_{j}$ are the components of the state vector. The time evolution is given by the Schrödinger equation
$\frac{d}{dt}|\psi\rangle=\frac{1}{ih}H|\psi\rangle,$ (2)
where $h$ is the Planck constant and $H$ is a $n\times n$ Hermitian matrix, the Hamiltonian matrix. This equation can also be written in terms of the components $z_{j}$. Replacing (1) into this equation we find
$\frac{dz_{j}}{dt}=\frac{1}{ih}\sum_{k}H_{jk}z_{k},$ (3)
where $H_{jk}$ are the elements of the matrix $H$, which holds the Hermitian property $H_{jk}^{*}=H_{kj}$.
If we discretize the time in equal interval of time $\tau$ and denoting by $|\psi^{\prime}\rangle$ the state vector at the instant $t+\tau$, then we may write
$|\psi^{\prime}\rangle=|\psi\rangle+\frac{\tau}{ih}H|\psi\rangle.$ (4)
Next we turn $z_{j}$ into stochastic variable that obey a stochastic equation which we establish by adding a noise into equation (4). We assume the following stochastic equation
$|\psi^{\prime}\rangle=e^{i\sqrt{\tau}\xi}|\psi\rangle+\frac{\tau}{ih}H|\psi\rangle,$ (5)
where $\xi$ is a random variable with zero mean and variance $\overline{\xi^{2}}=\gamma$. It is clear from this equation that the noise changes the phase of the state vector $|\psi\rangle$ but not its absolute value.
From equation (5), we find
$\langle\psi^{\prime}|\psi^{\prime}\rangle-\langle\psi|\psi\rangle=o(\tau),$ (6)
where $o(\tau)$ stands for terms of order strictly lower than $\tau$. Dividing by $\tau$ and taking the limit $\tau\to 0$, we find
$\frac{d}{dt}\langle\psi|\psi\rangle=0,$ (7)
which means that the norm is strictly conserved along the stochastic trajectory of the state vector in the Hilbert space. From now on we set the norm equal to one, that is,
$\langle\psi|\psi\rangle=\sum_{j}z_{j}z_{j}^{*}=1.$ (8)
The equation (5) can be written in terms of the components $z_{j}$ of the state vector. If we let $z_{j}^{\prime}$ be the component $z_{j}$ at the instant $t+\tau$ then
$z_{j}^{\prime}=e^{i\sqrt{\tau}\xi}z_{j}+\frac{\tau}{ih}\sum_{k}H_{jk}z_{k}.$ (9)
Up to terms of order $\tau$ this equation can be written as
$z_{j}^{\prime}=(1+i\sqrt{\tau}\xi-\frac{\tau}{2}\xi^{2})z_{j}+\frac{\tau}{ih}\sum_{k}H_{jk}z_{k}.$ (10)
Defining the increment $\Delta z_{j}=z_{j}^{\prime}-z_{j}$, we may write
$\Delta z_{j}=(i\sqrt{\tau}\xi-\frac{\tau}{2}\xi^{2})z_{j}+\frac{\tau}{ih}\sum_{k}H_{jk}z_{k}.$ (11)
III Underlying classical system
We show here that the equations (3) are equivalent to the Hamilton equations of motion of a classical system of $n$ degrees of freedom, which we call the underlying system. We denote by $q_{j}$ and $p_{j}$ the coordinate and the momentum associated to a degree of freedom. They form a pair of canonically conjugate variables, and the Hamilton equations of motion are
$\frac{dq_{j}}{dt}=\frac{\partial{\cal H}}{\partial p_{j}},\qquad\frac{dp_{j}}{dt}=-\frac{\partial{\cal H}}{\partial q_{j}}$ (12)
[…]
$z_{j}=\alpha q_{j}+i\beta p_{j}\qquad z_{j}^{*}=\alpha q_{j}-i\beta p_{j}$ (13)
where $z_{j}$ and its complex conjugate $z_{j}^{*}$ are understood as independent variables, and $\alpha$ and $\beta$ are constant to be chosen. Using this transformation, and choosing $\alpha$ and $\beta$ such that $2\alpha\beta=1/h$, we find the equations
$\frac{dz_{j}}{dt}=\frac{1}{ih}\frac{\partial{\cal H}}{\partial z_{j}^{*}}\qquad\frac{dz_{j}^{*}}{dt}=-\frac{1}{ih}\frac{\partial{\cal H}}{\partial z_{j}}$ (14)
[…]
${\cal H}=\sum_{jk}H_{jk}z_{k}z_{j}^{*}$ (15)
Replacing this form in the first of the equations (14) we reach (3). If we replace it in the second of the equations (14) we reach an equation which is the complex conjugate of (3).
The introduction of the underlying system allows to interpret the real part and the imaginary parts of the component $z_{j}$ of the state vector as proportional to the coordinate $q_{j}$ and momentum $p_{j}$ associated to a degree of freedom. We remark that the underlying system is not to be confused or identified with the real quantum system and in this sense it could be called fictitious. But then we should also call the wave function fictitious. Therefore, it is more appropriate to call the underlying system an unobservable as much as the wave function is an unobservable.
IV Fokker-Planck equation
From the stochastic equation of motion we may determine an equation for the time evolution of the probability distribution ${\cal P}(z,z^{*})$ of the set of variables $z_{j}$ and $z_{j}^{*}$. To this end we start by considering a bilinear function
${\cal F}=\sum_{jk}F_{jk}z_{j}z_{k}^{*}.$ (16)
The variation of this function between to instants of time $t+\tau$ and $t$ is given by
$\Delta{\cal F}=\sum_{j}(\frac{\partial{\cal F}}{\partial z_{j}}\Delta z_{j}+\frac{\partial{\cal F}}{\partial z_{j}^{*}}\Delta z_{j}^{*})$
$\sum_{jk}\frac{\partial^{2}{\cal F}}{\partial z_{j}\partial z_{k}^{*}}\Delta z_{j}\Delta z_{k}^{*}.$ (17)
Using (10) and the abbreviation
$f_{j}=\frac{1}{ih}\sum_{k}H_{jk}z_{k},$ (18)
then we find up to terms of order $\tau$
$\Delta{\cal F}=\sum_{j}\frac{\partial{\cal F}}{\partial z_{j}}(\tau f_{j}-\frac{\sqrt{\tau}}{2}\xi z_{j}-\frac{\tau}{2}\xi^{2}z_{j})+$
$\sum_{j}\frac{\partial{\cal F}}{\partial z_{j}^{*}}(\tau f_{j}^{*}-\frac{\sqrt{\tau}}{2}\xi z_{j}^{*}-\frac{\tau}{2}\xi^{2}z_{j}^{*})+$
$\tau\xi^{2}\sum_{jk}\frac{\partial^{2}{\cal F}}{\partial z_{j}\partial z_{k}^{*}}z_{j}z_{k}^{*}.$ (19)
From which follows
$\frac{d\langle{\cal F}\rangle}{dt}=\sum_{j}\langle f_{j}\frac{\partial{\cal F}}{\partial z_{j}}+f_{j}^{*}\frac{\partial{\cal F}}{\partial z_{j}^{*}}\rangle$
$\frac{\gamma}{2}\sum_{j}\langle z_{j}\frac{\partial{\cal F}}{\partial z_{j}}+z_{j}^{*}\frac{\partial{\cal F}}{\partial z_{j}^{*}}\rangle+\gamma\sum_{jk}\langle\frac{\partial^{2}{\cal F}}{\partial z_{j}\partial z_{k}^{*}}z_{j}z_{k}^{*}\rangle,$ (20)
where the averages are obtained using the probability distribution ${\cal P}(z,z^{*})$.
The equation that gives the time evolution of the probability distribution is obtained from (20) and is given by
$\frac{\partial\cal P}{\partial t}=-\sum_{j}(\frac{\partial f_{j}{\cal P}}{\partial z_{j}}+\frac{\partial f_{j}^{*}{\cal P}}{\partial z_{j}^{*}})$
$\frac{\gamma}{2}\sum_{j}(\frac{\partial z_{j}{\cal P}}{\partial z_{j}}+\frac{\partial z_{j}^{*}{\cal P}}{\partial z_{j}^{*}})+\gamma\sum_{jk}\frac{\partial^{2}z_{j}z_{k}^{*}{\cal P}}{\partial z_{j}\partial z_{k}^{*}}.$ (21)
which is a Fokker-Planck equation vankampen1981 ; tome2015 .
V Covariances
From the probability distribution $\cal P$}, we may determine $\langle z_{j}\rangle$, and the covariances $\langle z_{j}z_{k}^{*}\rangle$. The time evolution of these quantities can be obtained from the equation (20) and they are given by
$\frac{d\langle z_{j}\rangle}{dt}=\langle f_{j}\rangle-\frac{\gamma}{2}\langle z_{j}\rangle,$ (22)
and
$\frac{d\langle z_{j}z_{k}^{*}\rangle}{dt}=\langle f_{j}z_{k}^{*}+f_{k}^{*}z_{j}\rangle,$ (23)
and we see that the terms involving $\gamma$ disappears from this last equation. Using (18) we obtain
$\frac{d\langle z_{j}z_{k}^{*}\rangle}{dt}=\frac{1}{ih}\sum_{l}(H_{jl}\langle z_{l}z_{k}^{*}\rangle-H_{lk}\langle z_{l}^{*}z_{j}\rangle).$ (24)
Defining the density matrix $\rho_{jk}$ by
$\rho_{jk}=\langle z_{j}z_{k}^{*}\rangle,$ (25)
we reach the equation
$\frac{d\rho_{jk}}{dt}=\frac{1}{ih}\sum_{l}(H_{jl}\rho_{lk}-H_{lk}\rho_{jl}),$ (26)
which is the quantum Liouville equation. We observe, however, that from the derivation of this equation that $\rho_{jk}$ is understood as the covariances of the components of the stochastic state vector $|\psi\rangle$. Since the norm is equal to one then
$\sum_{j}\rho_{jj}=1.$ (27)
Defining the $n\times n$ matrix $\rho$ whose elements are $\rho_{jk}$ then the equation (26) can be written in the matrix form
$\frac{d\rho}{dt}=\frac{1}{ih}(H\rho-\rho H).$ (28)
The Liouville equation has a special type of solution of the type $\rho_{jk}(t)=\phi_{j}(t)\phi_{k}^{*}(t)$. Indeed, if we replace this form in equation (26) we find
$\phi_{j}\frac{d\phi_{k}^{*}}{dt}+\phi_{k}^{*}\frac{d\phi_{j}}{dt}=\frac{\phi_{k}^{*}}{ih}\sum_{l}H_{jl}\phi_{l}-\frac{\phi_{j}}{ih}\sum_{l}H_{lk}\phi_{l}^{*},$ (29)
and we see that this is possible if
$\frac{d\phi_{j}}{dt}=\frac{1}{ih}\sum_{l}H_{jl}\phi_{l},$ (30)
and as long as the initial condition is such that $\rho_{jk}(0)=\phi_{j}(0)\phi_{k}^{*}(0)$ for all pairs $jk$.
If we define the vector
$|\psi_{s}\rangle=\sum_{j}\phi_{j}|j\rangle.$ (31)
Then using (30) it follows that it obeys the equation
$\frac{d}{dt}|\psi_{s}\rangle=\frac{1}{ih}H|\psi_{s}\rangle,$ (32)
which is identified with the Schrödinger equation. We call $|\psi_{s}\rangle$ the Schrodinger state vector which should not be confused with the stochastic state vector $|\psi\rangle$ given by (1) which obey the stochastic equation (5).
It is worthwhile to compare the equations obeyed by these two state vector. To this end we write (32) in the discrete time form
$|\psi_{s}^{\prime}\rangle=|\psi_{s}\rangle+\frac{\tau}{ih}H|\psi_{s}\rangle,$ (33)
which should be compare with (5) which we rewrite here
$|\psi^{\prime}\rangle=e^{i\sqrt{\tau}\xi}|\psi\rangle+\frac{\tau}{ih}H|\psi\rangle,$ (34)
and we recall that $\xi$ is a random variable, absent in the previous equation.
The relation between the components of the the Schrödinger state vector and the stochastic state vector is
$\phi_{j}\phi_{k}^{*}=\langle z_{j}z_{k}^{*}\rangle,$ (35)
and in particular
$|\phi_{j}|^{2}=\langle|z_{j}|^{2}\rangle,$ (36)
and, according to the present approach, $|\phi_{j}|^{2}$ is not properly a probability but is a variance associated to the stochastic state vector, and the off diagonal terms $\phi_{j}\phi_{k}^{*}$ are the covariances associated to the stochastic state vector.
VI Continuous space
We now consider an infinite-dimensional Hilbert space. We assume that the wave function $\psi(x,t)$ is given by
$\psi(x)=\sum_{j}c_{j}\phi_{j}(x),$ (37)
where $\phi_{j}(x)$ are the eigenfunctions of a certain diferential operator which forms a complete set of orthonormalized functions. The $c_{j}$ are coeffients which depend on time and are understood as stochastic variables.
In an analogous way we assume the following stochastic equation in discrete time
$\psi^{\prime}=e^{i\sqrt{\tau}\xi}\psi+\frac{\tau}{ih}\hat{H}\psi,$ (38)
where here $\hat{H}$ is the Hamiltonian differentical operator and $\psi^{\prime}$ is the wave function at time $t+\tau$. Using the same reasoning as before, we conclude that the norm of the stochastic wave function is preserved at any stochastic trajectory,
$\frac{d}{dt}\int|\psi(x)|^{2}=0,$ (39)
which allows to set it equal to a constant which we choose to be one, that is,
$\int|\psi(x)|^{2}=1.$ (40)
Replacing (37) in (38), we find
$c_{j}^{\prime}=e^{i\sqrt{\tau}\xi}c_{j}+\frac{\tau}{ih}\sum_{k}H_{jk}c_{k},$ (41)
where
$H_{jk}=\int\phi_{j}^{*}\hat{H}\phi_{k}dx,$ (42)
and $c_{j}^{\prime}$ stands for $c_{j}$ at time $t+\tau$.
Equation (41) becomes identical to (10) if we formally replace $c_{j}$ by $z_{j}$. Therefore we may draw the same conclusions. For instante, the equation (26) becomes
$\frac{d\rho_{jk}}{dt}=\frac{1}{ih}\sum_{l}(H_{jl}\rho_{lk}-H_{lk}\rho_{jl}),$ (43)
where here
$\rho_{jk}=\langle c_{j}c_{k}^{*}\rangle,$ (44)
Defining
$\rho(x,x^{\prime})=\langle\psi(x)\psi^{*}(x^{\prime})\rangle,$ (45)
then
$\rho(x,x^{\prime})=\sum_{jk}\rho_{jk}\phi_{j}(x)\phi_{k}^{*}(x^{\prime}).$ (46)
Deriving with respect to time and using (43) we find
$\frac{\partial}{\partial t}\rho(x,x^{\prime})=\frac{1}{ih}\int H(x,x^{\prime\prime})\rho(x^{\prime\prime},x^{\prime})dx^{\prime\prime}$
$\frac{1}{ih}\int\rho(x,x^{\prime\prime})H(x^{\prime\prime},x^{\prime})dx^{\prime\prime},$ (47)
where
$H(x,x^{\prime})=\sum_{jk}\phi_{j}(x)H_{jk}\phi_{k}^{*}(x^{\prime}),$ (48)
$\rho(x,x^{\prime})=\sum_{jk}\phi_{j}(x)\rho_{jk}\phi_{k}^{*}(x^{\prime}).$ (49)
Equation (47) can also be written in the form
$\frac{\partial}{\partial t}\rho(x,x^{\prime})=\frac{1}{ih}[\hat{H}_{x}\rho(x,x^{\prime})-\hat{H}_{x^{\prime}}\rho(x,x^{\prime})],$ (50)
where the differential operators $\hat{H}_{x}$ and $\hat{H}_{x^{\prime}}$ act on functions of $x$ and $x^{\prime}$, respectively.
Let us consider solutions of the type
$\rho(x,x^{\prime})=\psi_{s}(x)\psi_{s}^{*}(x^{\prime}).$ (51)
If we replace in (50), we see that this is indeed a solution as long as
$\frac{\partial\psi_{s}(x)}{\partial t}=\frac{1}{ih}\hat{H}_{x}\psi_{s}(x),$ (52)
which identify with the Schrödinger equation. Notice that $\psi_{s}(x)$ should no be confused with $\psi(x)$ given by (37) which we call stochastic wave function because the coefficients $c_{j}$ are stochastic variables. The solutions of the type (51) allow us to write $\rho_{jk}=\langle c_{j}c_{k}^{\rangle}=a_{j}a_{k}^{*}$, in which case the Schrödinger wave function is written as
$\psi_{s}(x)=\sum_{j}a_{j}\phi_{j}(x).$ (53)
It is worth comparing the equations obeyed by $\psi$ and by $\psi_{s}$. The equation obeyed by the former is the equation (38), which we rewrite here
$\psi^{\prime}=e^{i\sqrt{\tau}\xi}\psi+\frac{\tau}{ih}\hat{H}\psi.$ (54)
The equation obeyed by $\psi_{s}$ is the equation (52) which in the discretize form is
$\psi_{s}^{\prime}=\psi_{s}+\frac{\tau}{ih}\hat{H}_{x}\psi_{s},$ (55)
where $\psi_{s}^{\prime}$ is the wave function at time $t+\tau$.
Let us look at the density $\rho(x,x)=|\psi_{s}(s)|^{2}$. In the standard interpretation, $|\psi_{s}(s)|^{2}dx$ is the probability of finding the particle between $x$ and $x+dx$. In the present approach $|\psi_{s}(s)|^{2}$ is the variance $\langle|\psi(x)|^{2}\rangle$ of the stochastic wave function $\psi(x)$.
VII Underlying continuous classical system
In the case of continuous space we can also introduce an underlying classical system whose equations of motion are equivalent to Schrödinger equation. In this case the Hamilton equations of motion are
$\frac{\partial\psi}{\partial t}=\frac{1}{ih}\frac{\delta{\cal H}}{\delta\psi^{*}},\qquad\frac{\partial\psi^{*}}{\partial t}=-\frac{1}{ih}\frac{\delta{\cal H}}{\delta\psi},$ (56)
[…]
${\cal H}=\int\psi^{*}(x)H(x,x^{\prime})\psi(x^{\prime})dxdx^{\prime},$ (57)
and $H(x,x^{\prime})$ is given by (48). Calculating the functional derivative we find
$\frac{\partial\psi(x)}{\partial t}=\frac{1}{ih}.\int H(x,x^{\prime})\psi(x^{\prime})dx^{\prime}$ (58)
Taking into account the relations (48) and (42), we find
$\int H(x,x^{\prime})\psi(x^{\prime})dx^{\prime}=\hat{H}\psi(x),$ (59)
and we may write
$\frac{\partial\psi}{\partial t}=\frac{1}{ih}\hat{H}\psi.$ (60)
If we replace (59) into (57), we get
${\cal H}=\int\psi^{*}(x)\hat{H}\psi(x)dx.$ (61)
[…]
$\langle{\cal H}\rangle=\int H(x,x^{\prime})\rho(x^{\prime},x)dxdx^{\prime}.$ (62)
If $\rho(x^{\prime},x)=\psi_{s}(x^{\prime})\psi_{s}(x)^{*}$ then
$\langle{\cal H}\rangle=\int\psi_{s}^{*}(x)H(x,x^{\prime})\psi_{s}(x^{\prime})dxdx^{\prime}.$ (63)
Which can also be written as
$\langle{\cal H}\rangle=\int\psi_{s}^{*}(x)\hat{H}\psi_{s}(x)dx.$ (64)
VIII Position and momentum
Let us write for the position and momentum equations analogous to (57) and (61). For the position we have
${\cal X}=\int\psi^{*}(x)\hat{x}\psi(x)dx,$ (65)
where $\hat{x}=x$, and
${\cal X}=\int\psi^{*}(x)X(x,x^{\prime})\psi(x^{\prime})dxdx^{\prime},$ (66)
where
$X(x,x^{\prime})=x\delta(x-x^{\prime})=\frac{x}{2\pi}\int e^{i(x-x^{\prime})k}dk,$ (67)
and
$\langle{\cal X}\rangle=\int X(x,x^{\prime})\rho(x^{\prime},x)dxdx^{\prime}.$ (68)
For the momentum, we have
${\cal Y}=\int\psi^{*}(x)\hat{p}\psi(x)dx,$ (69)
where $\hat{p}=-i\hbar\partial/\partial x$, and
${\cal Y}=\int\psi^{*}(x)Y(x,x^{\prime})\psi(x^{\prime})dxdx^{\prime},$ (70)
where
$Y(x,x^{\prime})=i\hbar\frac{1}{2\pi}\int ke^{i(x-x^{\prime})k}dk,$ (71)
and
$\langle{\cal Y}\rangle=\int Y(x,x^{\prime})\rho(x^{\prime},x)dxdx^{\prime}.$ (72)
The averages $\langle{\cal X}\rangle$ and $\langle{\cal Y}\rangle$ are understood as the values obtained in the measurement of the position and momentum of a particle. According to formulas (68) and (70), they are related to the covariances of the stochastic wave function. Let us compare with the standart interpretation. To this end let us write $\cal X$} as
${\cal X}=\int\psi^{*}(x)x\psi(x)dx,$ (73)
and
$\langle{\cal X}\rangle=\int x\rho(x,x)dx.$ (74)
In the standard interpretation, $\rho(x,x)dx$ is the probability of finding the particle between $x$ and $x+dx$, that is, $\rho(x,x)$ is probability density distribution associated to the variable $x$, which is thus understood as a random variable. Here, on the other hand, $\rho(x,x)$ is the variance of $\psi(x)$, understood as a stochastic variable, associated to a certain probability distribution. We did not write down the probability distribution for the present case of continuous variable. But it is similar to that of discrete variables, which obeys the Fokker-Planck equation (21).
IX Conclusion
The probabilistic character of quantum mechanics was expressed here by the introduction of a stochastic dynamics for the wave function. This was implemented by assuming that the wave function is a stochastic variable that obeys a stochastic equation. From this equation we show that the density matrix is a covariance matrix. That is, the elements of the density matrix are the covariances of the stochastic wave functions. We used a noise that changes the phase of the wave function but not its absolute value. This implies that the stochastic dynamics preserves strictly the norm of the wave function.
We have also shown that a quantum system defined on a Hilbert space of dimension $n$ can by described by an underlying classical system of $n$ degrees of freedom. The real and imaginary parts of the wave function are proportional to the coordinate and the momentum associated to a certain degree of freedom.
From our results, the wave function function follows a stochastic trajectory in phase space so that at each time we may associate a probabilistic character of the quantum system and not only at the instant of measurement. Thus according to the present approach, the collapse of the wave function is not required.

### v2

@@
 However, we need also to state the meaning or the interpretation of probability.
-Within the standard interpretation, it seems that it is interpretated as related to frequency, which can be measured experimentally.
+Within the standard interpretation, it seems that it is interpreted as related to frequency, which can be measured experimentally.
 In accordance with the standard interpretation, the probabilistic character of the wave function emerges at the instant of measurement, when the wave function is said to collapse.
@@
 To answer this question we must take a measurement at the instant $t^{\prime}$.
-But, if we do that the wave function collapses at $t$ and the measurement at time $t$ becomes compromised.
-From these reasonings, we may say that the probabilistic character of the wave function prior to measurement is not li ble to measurement.
+But, if we do that the wave function collapses at $t^{\prime}$ and the measurement at time $t$ becomes compromised.
+From these reasonings, we may say that the probabilistic character of the wave function prior to measurement is not liable to measurement.
 Within the standard interpretation of quantum mechanics, it is assumed that the wave function is not associated to probability prior to measurement, and that the probabilistic character emerge at the instant of measurement.
 Our aim here is to provide an approach in which the wave function does have a probabilistic character at any time.
-This is acomplished by treating the wave function as a stochastic variable which obeys a stochastic equation of motion oliveira2024a ; oliveira2024b ; oliveira2025 .
+This is accomplished by treating the wave function as a stochastic variable which obeys a stochastic equation of motion oliveira2024a ; oliveira2024b ; oliveira2025 .
 It is important to say that the probability character at the instant of measurement given by the present approach is the same as that given by the standard interpretation.
+Our stochastic approach should not be confused or identified with previous stochastic approaches nelson1966 in which the stochastic variable is the quantum position of the particle and not the wave function.
+Notice that in our approach the number of stochastic variable is proportional to the dimension of the Hilbert space whereas in these previous stochastic approaches it is equal to the number of degrees of freedom of the quantum system, which is one for a quantum particle moving on a straight line.
+From our approach it follows that the covariance matrix associated to the stochastic wave function is identified as the quantum density matrix.
+The equation that one obtains for the covariances, which is a closed equation in the covariances, is then identified with the quantum Liouville equation.
+This equation may have solutions consisting of the product of a function and its complex conjugate.
+When this happens these functions are identified as the pure states.
+The stochastic equations that we use are the classical Hamiltonian equations of motion complemented by a noise.
+The noise we use is a special type of noise that preserves the norm of the stochastic wave function.
+This is not merely a technical detail but has a conceptual significance because without it we would not obtain quantization properties from classical motion.
+We are interested here only in the part of the interpretation of quantum mechanics associated with the probabilistic aspect considering our stochastic treatment of the wave function.
+This treatment allows to understand that the probabilistic character of quantum mechanics occurs at all times, and not just at the time of measurement, thus making the collapse of the wave function unnecessary.
 II Stochastic equations of motion
@@
 $\psi(x)=\sum_{j}c_{j}\phi_{j}(x),$ (37)
-where $\phi_{j}(x)$ are the eigenfunctions of a certain diferential operator which forms a complete set of orthonormalized functions.
-The $c_{j}$ are coeffients which depend on time and are understood as stochastic variables.
+where $\phi_{j}(x)$ are the eigenfunctions of a certain differential operator which forms a complete set of orthonormalized functions.
+The $c_{j}$ are coefficients which depend on time and are understood as stochastic variables.
 In an analogous way we assume the following stochastic equation in discrete time
 $\psi^{\prime}=e^{i\sqrt{\tau}\xi}\psi+\frac{\tau}{ih}\hat{H}\psi,$ (38)
-where here $\hat{H}$ is the Hamiltonian differentical operator and $\psi^{\prime}$ is the wave function at time $t+\tau$.
+where here $\hat{H}$ is the Hamiltonian differential operator and $\psi^{\prime}$ is the wave function at time $t+\tau$.
 Using the same reasoning as before, we conclude that the norm of the stochastic wave function is preserved at any stochastic trajectory,
@@
 and $c_{j}^{\prime}$ stands for $c_{j}$ at time $t+\tau$.
-Equation (41) becomes identical to (10) if we formally replace $c_{j}$ by $z_{j}$. Therefore we may draw the same conclusions. For instante, the equation (26) becomes
+Equation (41) becomes identical to (10) if we formally replace $c_{j}$ by $z_{j}$. Therefore we may draw the same conclusions. For instance, the equation (26) becomes
 $\frac{d\rho_{jk}}{dt}=\frac{1}{ih}\sum_{l}(H_{jl}\rho_{lk}-H_{lk}\rho_{jl}),$ (43)
@@
 Notice that $\psi_{s}(x)$ should no be confused with $\psi(x)$ given by (37) which we call stochastic wave function because the coefficients $c_{j}$ are stochastic variables.
-The solutions of the type (51) allow us to write $\rho_{jk}=\langle c_{j}c_{k}^{\rangle}=a_{j}a_{k}^{*}$, in which case the Schrödinger wave function is written as
+The solutions of the type (51) allow us to write $\rho_{jk}=\langle c_{j}c_{k}^{*}\rangle=a_{j}a_{k}^{*}$, in which case the Schrödinger wave function is written as
 $\psi_{s}(x)=\sum_{j}a_{j}\phi_{j}(x).$ (53)